# CADDesigner: Conceptual Design of CAD Models Based on General-Purpose Agent

### v2

@@ Introduction
We compared our method with two learning based text-to-CAD generation methods: Text2CAD (Khan et al. 2024b),cadrille (Kolodiazhnyi et al. 2025) and one agent-based method: CADCodeVerify (Wu et al. 2023a).
@@
  • cadrille proposes a two-stage CAD reconstruction pipeline: supervised fine-tuning followed by reinforcement learning.
@@
We sampled $1k$ instances in an even distribution of model complexity from the $8k$ test set of Text2CAD for comparison. For each instance, we used the abstract text description as input prompts. We then performed inference using their released weights and our method to generate results for metric calculation. As presented in Table 2 and Figure 3, the experimental results show that the model generated by our method achieves the optimal performance. Our method demonstrates the best prompt-result alignment, followed by CADCodeVerify and Text2CAD. Cadrille fails to generate valid outputs due to poor generalization, as it was trained on expert-level descriptions and cannot adapt to abstract-level inputs during inference.
Table 2: Performance comparison of different Text-to-CAD methods under abstract text inputs.
@@
Limitations
Although our method demonstrates considerable flexibility, it still faces challenges in generating models that require precise geometric constraints. While input text or images can explicitly specify numerical parameters (e.g., dimensions), current large language models (LLMs) and vision-language models (VLMs) exhibit limited understanding and reasoning abilities of spatial geometric relationships and topological dependencies. This often results in outputs that satisfy local constraints (e.g., correct height or diameter of a cup) but violate global consistency (e.g., cup’s handle is inside the cup).
Due to limitations in the current code-generation models, the approach underperforms in generating CAD models involving complex mathematical computations, such as involute gears. These tasks, requiring calculations of fundamental geometric parameters, meshing relationships, or strength validation, yet cannot perform out well due to the complexity of generating correct spatial geometry calculation code.
Conclusion and Future Work
We propose a novel framework CADDesigner for CAD modeling code generation combined with a context-independent CAD modeling script generation paradigm to generate high-quality CAD modeling scripts. Experiments show that this method can achieve SOTA results in the field of CAD conceptual design.
CADDesigner is highly effective for rapidly prototyping CAD models, capable of generating innovative CAD designs and producing high-quality, complex datasets that align text descriptions with corresponding CAD modeling code. Future work could integrate learning-based methods to accept point clouds and B-rep data as inputs, enabling the model to learn geometric and topological constraints through training. Additionally, incorporating models specialized in spatial geometry reasoning could assist CADDesigner in computing complex CAD parameters, such as those required for gear and crankshaft design.

### v4

@@ 5.5 Comparison Methods
We compared our method with two learning-based text-to-CAD generation methods: Text2CAD [text2cad], cadrille [kolodiazhnyi2025cadrille] and one agent-based method: CADCodeVerify [wu2023visual].
@@
  • cadrille proposes a two-stage CAD reconstruction pipeline: supervised fine-tuning (SFT) on large-scale procedurally generated data, followed by reinforcement learning fine-tuning using online feedback.
@@
For evaluation, we use 1k samples from the filtered test set described in Section 5.2. For each instance, we used the abstract text description as input prompts. We then performed inference using their released weights and our method to generate results for metric calculation. As presented in Table 3 and Figure 4, the experimental results show that the model generated by our method achieves the optimal performance. Our method demonstrates the best prompt-result alignment, followed by CADCodeVerify and Text2CAD. While cadrille achieves a high SUC score, indicating it can produce syntactically valid CAD models in most cases, the visualized results reveal these outputs are typically meaningless with respect to the input instructions (see Figure 4). This failure is due to cadrille’s poor generalization from expert-level to abstract instructions: it generates structurally valid but semantically irrelevant models. In contrast, the other methods generate outputs that are both syntactically and semantically valid to varying degrees, with our method being the most consistent.
Table 3: Performance comparison of different Text-to-CAD methods under abstract text inputs on 1K test models.
@@
Figure 4: Comparison of generation results across Text2CAD, CADCodeVerify, cadrille, and our method. CADDesigner achieves the best input alignment. Text2CAD and CADCodeVerify show moderate performance, while cadrille generates syntactically valid but semantically incorrect outputs due to poor generalization from expert-level training to abstract inputs.
Figure 5: Comparison of CADDesigner and cadrille on image-based inputs. CADDesigner (blue) generates structurally accurate CAD models that closely match the input images by supporting expressive operations such as revolve and high-level patterns, whereas cadrille (green), relying on low-level extrusion, often produces syntactically valid but geometrically less faithful results.
Additionally, we compare our method with cadrille on image-based inputs to further illustrate their respective modeling capabilities. As shown in Figure 5, CADDesigner consistently generates CAD models that closely match the input images, benefiting from a more expressive CAD operation space. In particular, several models require rotationally symmetric structures that are naturally constructed via revolve operations, which are explicitly supported by CADDesigner but not by cadrille, leading to missing or distorted geometries in the latter. Moreover, models with rich geometric details and repetitive structures are more effectively specified using high-level CAD operations such as arrays or patterned constructs. CADDesigner can leverage these higher-level APIs to compactly and precisely generate such structures, whereas cadrille relies primarily on low-level extrusion-based operations, often resulting in geometrically less faithful outputs.
These results highlight that beyond syntactic validity, the availability of expressive and semantically aligned CAD operations is crucial for faithful image-based CAD model generation.
5.6 Limitations
Although our method demonstrates considerable flexibility, it still faces challenges in generating models that require precise geometric constraints. While input text or images can explicitly specify numerical parameters (e.g., dimensions), current LLMs and VLMs exhibit limited understanding and reasoning abilities of spatial geometric relationships and topological dependencies. This often results in outputs that satisfy local constraints (e.g., correct height or diameter of a cup) but violate global consistency (e.g., the cup’s handle is inside the cup).
Due to limitations in the current code-generation models, the approach underperforms in generating CAD models involving complex mathematical computations, such as involute gears. These tasks, requiring calculations of fundamental geometric parameters, meshing relationships, or strength validation, cannot be performed well due to the complexity of generating correct spatial geometry calculation code.
6 Conclusion and Future Work
We propose CADDesigner, a novel framework for CAD modeling code generation, combined with an explicit context imperative paradigm to produce high-quality CAD modeling scripts. Experimental results demonstrate that CADDesigner achieves state-of-the-art performance in the domain of conceptual CAD design. CADDesigner is particularly effective for rapid prototyping, enabling the generation of innovative CAD models and the construction of complex, high-quality datasets that align textual descriptions with corresponding CAD modeling code.
In future work, we plan to extend CADDesigner by incorporating learning-based methods that accept point clouds and B-rep data as input, allowing the system to learn geometric and topological constraints through data-driven training. Furthermore, integrating models specialized in spatial geometric reasoning may enhance CADDesigner’s ability to compute intricate CAD parameters, such as those required for designing gears, crankshafts, and other mechanically constrained components.